OpenStudy (anonymous):

solve system of equation y=5x-2 y=2x+10 AND x=-2y-1 2x+y=-20

OpenStudy (anonymous):

system 1: equate the y's, but write in terms of x: 5x-2=2x+10 3x=12 x=4 back sub for y y=5(4)-2=18

OpenStudy (anonymous):

make the first 2 equations equal to one another and solve for x, then for the last 2 equations isolate y for each side and do the same as before

OpenStudy (anonymous):

system 2: sub x from equation 1 into eqn 2, but write in terms of y 2(-2y-1)+y=-20 -4y-2+y=-20 -3y=-18 y=6 back sub into the first eqn to get x x=-2(6)-1=-12-1=-13